# Fairness in AI-Driven Recruitment: Challenges, Metrics, Methods, and Future Directions

### v2

@@ 40 Black 12 12/40 30%
TABLE I: Example selection rate disparity showing adverse impact.
The four-fifths rule has recently come under scrutiny, with proposed legislation advocating for a smaller disparity in selection rates [68]. From this and other legislation, two key definitions of discrimination have emerged in AI literature: disparate treatment, which refers to intentionally discriminatory practices targeted at individuals based on their protected attributes; and disparate impact, which encompasses practices that result in a disproportionately adverse effect on protected groups [67, 69, 70]. Although it might seem straightforward to meet these standards by simply removing protected attributes from the training data, many situations still arise where, despite these efforts—and sometimes because of them—unfair decisions are made. For instance, in natural language processing (NLP), a model might infer a candidate’s gender from their name, a phenomenon known as indirect discrimination or discrimination through features implicit in the dataset [69]. Therefore, to achieve fairness in a decision-making system, it is essential to address the underlying causes of bias in the data. This is discussed next.
II-C Causes of Bias
@@
    Proxies: Even with protected attributes removed from the training data, group membership can be inferred from other attributes, potentially leading to biased decisions. For instance, Amazon’s hiring application was biased even without using the gender attribute because the model could infer gender from the educational institution listed on the resumes of applicants (e.g., all-female or all-male colleges) [13].
@@
III Bias in AI-Based Recruitment
HR and I-O psychology literature typically divides the recruitment pipeline of an organization into four stages, as depicted in Figure 2 [21, 38, 75]. This section provides an overview of AI-driven practices, potential biases, and efforts to achieve fairness at each stage of the recruitment process.
Figure 2: Overview of steps in a typical recruitment pipeline: (1) candidate sourcing (after job analysis), (2) candidate screening, (3) candidate interviews, and (4) selection and offer, including post-decision evaluation.
III-A Sourcing
The first stage of the recruitment process begins with a job analysis, where an employer identifies a position within the organization and its required experience, knowledge, skills, abilities, and other pertinent characteristics (KSAOs). This stage also involves evaluating criteria for person-environment fit [72, 76, 77], which includes assessing how well an individual’s values and traits align with the company’s work culture.
@@ III-A Sourcing
In this context, AI-based methods have been developed for both job analysis and candidate sourcing stages. AI-based candidate sourcing platforms, such as HireEZ [78] and Arya [79], utilize AI to reach a broader and more diverse range of candidates [80]. Additionally, LLMs are employed for various tasks, including crafting job ads [81], conducting job analysis, and sourcing candidates through prompting-based interfaces like TalentGPT [82, 83]. Other AI-powered services, such as Textio, BrightHire, and SeekOut [84, 82, 85, 86], are also available to streamline the candidate sourcing process.
However, two prominent areas of bias can occur in AI-based sourcing software. The first area of bias is in the text of job ads, which have been found to use gendered keywords and phrases, as well as perpetuate age and racial biases [81, 87, 88, 89, 86]. To address this issue, services like Textio use AI to provide neutral language and reduce potential bias [81]. Unfortunately, many job ads are not tested for bias in their wording, which can negatively impact not only the time and resources spent by organizations to find qualified candidates but also the trust and self-efficacy, or confidence, of the jobseeker [90, 6, 4]. Biased job ads can also influence AI-driven job-ad recommendations, leading to skewed ads being shown to users, as observed in previous studies on Facebook and Google job ad recommendations [14, 12].
@@
Job recommendation systems, used by jobseekers to browse job ads, heavily rely on AI to match resumes with ads in their database, often leading to biased recommendations. Examples of such platforms include LinkedIn [91] and Indeed [92], where job postings are aggregated for jobseekers to search, while recruiters frequently search for passive candidates—individuals who are not actively seeking a job but fit a job ad well [38]. The recommendation algorithms on these sites typically use collaborative filtering [93], content-based filtering [94, 93], or multi-objective ranking approaches [95]. However, job ad content can lead to biased recommendations, as an estimated 60% - 70% of job ads posted online disproportionately represent industries for high-skilled workers and STEM occupations [96]. Moreover, ad delivery systems can be skewed along gender and racial lines, as evidenced by Facebook’s job recommendation engine in 2019 [14]. Additionally, gender inequality was observed on job boards such as Indeed, Monster, and CareerBuilder in 2018, with 12 of the 35 studied job titles showing significant group unfairness [97].
@@
To address such biases, various approaches have been developed. For example, Imana et al. distinguish between skew in ad delivery caused by protected attributes and skew due to differences in candidate qualifications [98]. Similarly, Chen et al. [99] tackle recency bias in AI-driven information retrieval, where recent content disproportionately influences model training, leading to bias. Additionally, Rus et al. [100] use a generative adversarial network to eliminate gender biases from word vector representations in job ad text and resumes. Meanwhile, Li et al. model the job recommendation task as a resource allocation problem, promoting fairness by maintaining consistency across users from different protected groups [101]. Despite these efforts to mitigate cognitive biases in job and candidate recommendations, such biases may persist through the training data [75].
III-B Candidate Screening
The next stage in recruitment, candidate screening (or candidate matching), involves filtering an initial pool of candidate profiles, such as resumes, based on person-job fit. During this process, large companies that handle thousands of applications often use ranking and screening tools. Examples of AI-based screening methods include those provided by HireVue [102], Plum [103], and SeekOut [82]. These services offer resume ranking and filtering systems, assessments of candidate job experience and knowledge, as well as personality evaluations [77, 38, 104, 72].
However, similar to other areas of AI use, screening models are susceptible to bias, as observed in past applications such as Amazon’s resume ranking application [13, 105]. Institutional and systemic biases are often reflected in resume screening and assessment models because the training data is based on previous hiring decisions, which may lack representation from different protected groups [75]. Additionally, these systems, often based on NLP, may generate harmful language, perpetuate stereotypes, or create unequal representation of different groups [106]. For instance, in 2016, popular word embeddings, which are frameworks used to represent text data as vectors for training in neural networks, were found to exhibit gender stereotypes due to training on Google News articles [106]. Later, GPT-3 was found to generate strong negative stereotypes against religious groups [34], and other LLMs like GPT-4 and BERT have also come under scrutiny for bias [107]. These biases can negatively impact both jobseekers and organizations. If less-qualified candidates are hired due to bias, it can lead to additional time and resources being required for training and onboarding, potentially resulting in lower employee retention and job satisfaction rates [90].
@@ III-B Candidate Screening
Recognizing the importance of screening processes, many researchers have focused on developing methods to detect and mitigate biases. For example, Pena et al. [108, 109] propose a testbed for resume screening algorithms to identify bias in candidate scoring across five scenarios with different protected attributes. Similarly, Deshpande et al. [110] tackle bias by reweighting tf-idf values based on fairness metrics when matching resumes with job descriptions. Burke et al. [111] introduce a mitigation method through candidate ranking. While these techniques represent an initial effort in bias mitigation, non-textual features within screening models, such as images (commonly found in online profiles), and key non-protected attributes like educational background may still contribute to perpetuating bias. This issue, where non-textual features such as visual and auditory cues contribute to bias, is particularly prevalent in the screening of candidate video or phone interviews.
III-C Interviewing Candidates
@@
The third stage in the recruitment process involves interviewing candidates to assess their KSAOs, work values, personality, and other traits in detail. Over the past decade, the recruitment process has increasingly relied on AI to evaluate these traits during interviews. Various AI modalities have been utilized, including chatbot systems for candidate interaction, automated video and phone interviews, and online question-answer assessments [102, 23, 32, 75, 23]. For instance, platforms like HireVue [102] and BrightHire [85] heavily depend on AI to evaluate video interviews, scoring candidates based on personality traits and their fit for the job.
However, in the realm of video- and image-driven AI applications, concerns have arisen regarding potential biases, as highlighted by various studies [112, 113, 114]. This issue became prominent in 2021 when HireVue ceased the use of its video-based candidate scoring systems, citing the absence of transparency in AI decision-making as a primary reason [32]. Additionally, in 2020, Illinois enacted the Artificial Intelligence Video Interview Act, the first regulation addressing AI in the interviewing of job candidates [115]. This law mandates that employers must inform applicants that their interview will be scored by an AI-driven system, provide the characteristics used by the AI system in scoring, and obtain consent from each applicant before the interview [115]. Furthermore, research into AI-based video interviewing systems and chatbots has revealed biases originating from the training data [116, 117]. For instance, investigations into the First Impressions dataset, used in the 2017 ChaLearn challenges at CVPR, uncovered biased pairwise labels assigned to participants’ personality traits based on gender and race [118, 119]. When these labels were used to construct a model, the biases were further amplified through the model’s predictions [118, 119].
To promote fair candidate screening, researchers have implemented various strategies to address these biases. For example, Yan et al. have proposed an approach using adversarial learning to mitigate bias according to fairness metrics in personality trait prediction [119]. Additionally, Booth et al. [120] have studied bias measurement in affective computing problems, particularly in the context of mock video interviews. Significant efforts have also been made to develop chatbot interviewing systems that ensure fairness. Researchers have explored the use of fairness-aware classifiers to mitigate harmful text content and identify biased prompts from users [121]. The survey by Xue et al. [121] provides a comprehensive overview of these approaches and potential biases in chatbot systems.
III-D Selection and Evaluation
@@
The final stage of the recruitment process involves selecting one or more job candidates, followed by negotiating an offer, which is then used to evaluate the organization’s overall recruitment process. AI has been incorporated in several ways to assist not only in selection but also in the negotiation of job offers. For example, AI is used during job offer negotiations to provide salary and training recommendations, thereby aiding in the onboarding process of newly hired employees [75]. This is exemplified by software like Oracle Recruiting Cloud, which not only forecasts the probability of a job candidate accepting an offer but also allows employers to experiment with various salary structures, benefits packages, and other incentives to enhance the likelihood of acceptance [75]. Additionally, Pactum AI [122] automates various negotiation processes, such as salary negotiations and contract discussions, leveraging a multitude of factors, including job market outlook.
@@
Although such AI seeks to create a fairer negotiation process by eliminating human biases, other forms of bias can still occur. For instance, a 2019 UpTurn report highlighted that Oracle and other AI-driven negotiation tools might reinforce asymmetric information between the job candidate and employer, potentially leading to an unfair process for the candidate [75]. Additionally, these tools may conflict with legislation prohibiting the use of salary history in hiring decisions and the Equal Pay Act of 1963, which mandates equal pay for individuals in the same position [75, 62].
@@
Job candidates who are not hired often face a lack of transparency regarding the AI-driven decision-making processes of employers. According to a survey conducted by The Talent Board, 53.5% of candidates reported receiving no feedback after the screening and interview stages, while 69.7% received no feedback after being turned down [123]. Although employers are not legally obligated to provide explanations [123], the absence of feedback can negatively impact the perceived fairness and trustworthiness of an AI-based recruitment process. Transparency is crucial at every stage of the process, including communicating the job outcome, as it enables candidates to learn and improve for future interviews. Previous studies have demonstrated that providing explanations for organizational decisions can significantly enhance trust [41, 38]. Recently, many approaches, as further described in Section IV-B, have prioritized model interpretability and explainability to enhance transparency in AI-based decision-making, a principle that can be extended to AI-driven hiring decisions.
@@
After finalizing hiring decisions, organizations evaluate their recruitment process based on three key criteria: validity, utility, and fairness. Validity ensures that assessments and interviews accurately reflect job KSAOs and predict job performance [38]. For instance, organizations often establish a criterion-referenced cut score as a cut-off for employee selection, based on the desired level of performance for a new employee [38]. Compliance with government guidelines, as described in Section II-B, is also essential to validity [38]. Next, utility assesses the cost-benefit ratio of staffing, which can be influenced by earlier recruitment stages such as employee outreach and job marketing. Finally, the fairness criteria is evaluated using employee selection guidelines described in Section II-A. However, as AI becomes more integrated into each stage of the process, it is necessary to incorporate bias detection and mitigation methods where guidelines may not adequately address algorithmic bias. For example, the uniform guidelines on employee selection could be substantially affected by the sample size of different groups. Work by Peng et al. [124] demonstrates how AI-human, or hybrid, decisions where AI assists humans, can still carry bias. Moreover, legislative definitions of fairness and employee selection guidelines may not be sufficiently detailed or up-to-date to address recent technological advancements.
IV Bias Detection and Mitigation

### v3

@@ 40 Black 12 12/40 30%
Table 1: Example illustrating selection rate disparity according to the EEOC’s four-fifths rule, demonstrating adverse impact.
However, the four-fifths rule and prior legislation are only the first step towards establishing guidelines to assess a fair hiring process using AI. For instance, the four-fifths rule has come under prior scrutiny with proposed legislation advocating for a smaller disparity in selection rates [87]. There have also been prior cases wherein these disparities have been contested on grounds of business necessity (that is, the criteria are directly job-related), cost considerations (alternatives are prohibitively expensive to implement), and qualification-based arguments (disparities result from differences in applicant qualifications rather than organizational practices) [88, 89].
Furthermore, although it might seem straightforward to meet these standards by simply removing protected attributes from the training data, many situations still arise where, despite these efforts—and sometimes because of them—unfair decisions are made. For instance, in natural language processing (NLP), a model might infer a candidate’s gender from their name, a phenomenon known as indirect discrimination or discrimination through features implicit in the dataset [85]. Similarly, legislation has often lagged behind AI advancements, with relevant laws only emerging in recent years and urging practitioners to act proactively. Therefore, achieving fairness in decision-making systems requires a nuanced approach that addresses the underlying causes of bias in the AI. This is discussed in the following section.
2.3 Causes of Bias
@@
    Proxies: Even with protected attributes removed from the training data, group membership can often be inferred from other attributes, potentially leading to biased decisions. For instance, word embeddings used to represent text in resumes have been shown to contain biased representations of gender, as evidenced by the distance between word vectors [94]. For example, although the term “computer programmer” is gender-neutral, its underlying vector representation may be closer to gender-specific terms such as “man,” causing it to serve as a proxy for gender [94]. However, in hiring, there are situations where proxies are necessary to verify credentials or to establish a bona fide occupational qualification (BFOQ)—for example, when a religious institution requires employees to be of a specific faith [95]. In such cases, protected attributes must be retained for fairness auditing, and potential bias mitigation techniques should be applied, as further described in Sec. 4.2.
@@
3 Bias in AI-Based Recruitment
Figure 2: Overview of steps in a typical AI-driven recruitment pipeline, including: (1) candidate sourcing (after job analysis), (2) candidate screening, (3) candidate interviews, and (4) selection and offer, with post-decision evaluation.
Hiring Stage Description References Summary of Key Challenges Most Relevant Fairness Metrics
Candidate Sourcing Forming a job description through job analysis, writing a job ad, and outreach to potential candidates via job boards or networking sites. [96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 11, 116, 14, 117, 118, 119, 120, 121, 122] Potential biased language in job postings discourages diverse applicants, perpetuated by AI tools trained on historically biased data, further limiting applicant diversity. Demographic parity, Counterfactual fairness, Multi-sided fairness
Candidate Screening Recruiters filter initial candidate pool by reviewing resumes and applications for job fit, typically without direct interaction. [123, 124, 99, 125, 126, 127, 128, 44, 45, 38, 94, 43, 129, 13, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139] AI models may prioritize quantifiable metrics over qualitative attributes, indirectly discriminate via proxies for protected attributes, and lack transparency in decision rationale. Accuracy parity, Demographic parity, Counterfactual fairness, Individual fairness, Predictive rate parity
Candidate Interviews Interaction with candidates to evaluate detailed KSAOs, personality traits, and cultural fit via video, phone, or chat-based AI systems. [123, 101, 36, 23, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 46, 159, 160] Biases embedded in large pre-trained models (LPTMs) from textual, audio, and visual data amplify implicit stereotypes, often resulting in accuracy disparities for non-native speakers or those with speech disabilities. Limited transparency and reliability concerns with automated personality assessments also increase regulatory scrutiny and fairness concerns. Individual fairness, Predictive rate parity, Accuracy parity, Procedural fairness, Counterfactual fairness, Fairness in regression
Selection & Offer Selecting final candidates, negotiating offers, and providing onboarding recommendations such as salary and benefits. [161, 162, 163, 164, 165, 166, 167] AI reliance on historical data perpetuates pay gaps, undervalues experience, penalizes resume gaps, and exacerbates information asymmetry during negotiation. Limited transparency in decision-making also negatively affects candidate perceptions of fairness. Individual fairness, Predictive rate parity, Procedural fairness, Fairness in regression
Table 2: Summary of surveyed research, fairness challenges, and most relevant fairness metrics identified at each stage of the AI-driven recruitment pipeline.
HR and I-O psychology literature typically divides the recruitment pipeline of an organization into four stages, as depicted in Figure 2 [21, 47, 168]. Each stage presents unique opportunities and challenges for bias introduction and mitigation. This section reviews current AI-driven practices at each stage, highlights potential sources of bias, and discusses existing efforts to achieve fairness throughout the recruitment process.
3.1 Candidate Sourcing
The first stage of the recruitment process begins with a job analysis, where an employer identifies a position within the organization and its required experience, knowledge, skills, abilities, and other pertinent characteristics (KSAOs). This stage also involves evaluating criteria for person-environment fit [91, 169], which includes assessing how well an individual’s values and traits align with the company’s work culture.
@@ 3.1 Candidate Sourcing
In this context, AI-based methods have been developed for both job analysis and candidate sourcing stages. AI-based candidate sourcing platforms, such as HireEZ [96], LinkedIn [111], and Arya [97], utilize AI to reach a broader and more diverse range of candidates. Other works, relying upon genetic algorithms [105] or large pre-trained models (LPTM) [106] have also been proposed. Additionally, LLMs are employed for various tasks, including crafting job ads [98], conducting job analysis [108, 109], and sourcing candidates through prompting-based interfaces like TalentGPT [99, 100]. They have also been used on the candidate side, by recommending jobs [107]. Other AI-powered services, such as Textio, BrightHire, Micro1, and SeekOut [102, 99, 101, 103, 110], are also available to streamline the candidate sourcing process. The advantage to AI usage in this stage is the efficient processing of candidate data, though a 2021 study noted a lack of trust in data accuracy, control, and transparency [104].
Bias in AI-based candidate sourcing generally originates from two prominent sources. First, bias arises in the textual content of job advertisements, which often include gendered keywords, phrases, and language perpetuating age or racial stereotypes [98, 122, 170, 171, 172, 103]. Such biased wording can deter certain demographic groups from applying, thereby limiting the diversity of the applicant pool. Services like Textio have sought to address this issue by employing AI-driven methods to suggest more neutral, inclusive language [98]. However, because many job advertisements remain untested for bias, this problem persists. Unchecked biases in job ads not only increase organizational time and resource expenditures in seeking diverse qualified candidates but also reduce the trust and self-efficacy—the confidence of jobseekers in their ability to succeed in the recruitment process—particularly among members of underrepresented or marginalized groups [5, 173]. Second, bias can stem from word embeddings in text-based models, stemming from inherent cultural, gender, racial, and political biases found in underlying training text data [94]. Biased job ads can also influence AI-driven job-ad recommendations, leading to skewed ads being shown to users, as observed in previous studies on Facebook and Google job ad recommendations [14, 11].
@@
Job recommendation systems, used by jobseekers to browse job ads, heavily rely on AI to match resumes with ads in their database, often leading to biased recommendations. Examples of such platforms include LinkedIn [111] and Indeed [112], where job postings are aggregated for jobseekers to search, while recruiters frequently search for passive candidates—individuals who are not actively seeking a job but fit a job ad well [47]. The recommendation algorithms on these sites typically use collaborative filtering [113], content-based filtering [114, 113], or multi-objective ranking approaches [115]. However, job ad content can lead to biased recommendations, as an estimated 60%–70% of job ads posted online disproportionately represent industries for high-skilled workers and STEM occupations [116]. Moreover, ad delivery systems can be skewed along gender and racial lines, as evidenced by Facebook’s job recommendation engine in 2019 [14]. Additionally, gender inequality was observed on job boards such as Indeed, Monster, and CareerBuilder in 2018, with 12 of the 35 studied job titles showing significant group unfairness [117].
@@
To address such biases, various approaches have been developed. For example, Imana et al. distinguish between skew in ad delivery caused by protected attributes and skew due to differences in candidate qualifications [118]. Similarly, Chen et al. [119] tackle recency bias in AI-driven information retrieval, where recent content disproportionately influences model training, leading to bias. Additionally, Rus et al. [120] use a generative adversarial network to eliminate gender biases from word vector representations in job ad text and resumes. Meanwhile, Li et al. model the job recommendation task as a resource allocation problem, promoting fairness by maintaining consistency across users from different protected groups [121]. Despite these efforts to mitigate cognitive biases in job and candidate recommendations, such biases may persist through the training data [168]. The issue of fairness in recommender systems extends beyond job recommendations and encompasses biases in e-commerce and content recommendation. It represents an active research area with ongoing efforts to address bias and enhance fairness [174].
3.2 Candidate Screening
The next stage, candidate screening, involves filtering a pool of candidate profiles (e.g., resumes and applications) based on their alignment with job requirements, or person-job fit [91, 47, 169]. To manage large volumes of applicants efficiently, organizations increasingly utilize AI-driven tools such as HireVue [123], Plum [124], and SeekOut [99]. These platforms offer automated resume ranking, candidate skill assessments, and preliminary evaluations of personality traits. Recent research has also explored advanced AI techniques, including traditional machine learning methods [133, 136], deep learning models like convolutional neural networks [135], recurrent neural networks [134], and LLMs [130, 131, 132].
However, similar to other areas of AI use, candidate screening models are susceptible to biases, as observed in high-profile cases such as Amazon’s AI-driven resume screening tool [13, 129]. Institutional biases inherent in historical hiring practices are frequently reflected in these models because their training data often lack diversity, perpetuating existing disparities [168]. Moreover, screening tools built upon NLP techniques are vulnerable to biases embedded within underlying training datasets, which can unintentionally reinforce harmful stereotypes or unequal representations of demographic groups [94]. For instance, widely used word embedding models were found to contain significant gender biases due to their training on biased textual data from sources like Google News [94]. Similarly, language models like GPT-3 have demonstrated biases against certain religious groups [38], and other prominent models, including GPT-4 and BERT, have faced scrutiny regarding biases related to gender and race [175]. Such biases in foundation models can directly lead to unfairness in resume screening and candidate selection processes, disadvantaging qualified candidates from underrepresented groups [44, 45, 43]. This bias not only undermines fairness but may also result in suboptimal hiring decisions, higher training costs, reduced productivity, and decreased employee retention and satisfaction, ultimately impacting organizational effectiveness and workplace diversity [176].
@@ 3.2 Candidate Screening
Recognizing the importance of screening processes, many researchers have focused on developing methods to detect and mitigate biases. Recent work by Wang et al. [137] have studied gender hiring bias in approaches using LLMs for resume scoring, and find seven out of ten studied LLMs show significant biases against males in at least one industry studied [137]. Similar studies on GPT-4o, Google Gemini 1.5, and Claude 3.5 have been conducted by An et al. [138], wherein they studied gender and racial bias in these models for evaluating resumes, finding lower resume scores were assigned for Black male candidates with similar work experience, education, and skills [138]. A similar study was conducted on LLM representations of occupations, and found perpetuation of gender stereotypes of jobs based on names; for instance, feminine first-name embeddings increased probabilities for female-dominated jobs [139]. Another approach by Pena et al. [125, 126] proposes a testbed for resume screening algorithms to identify bias in candidate scoring across five scenarios with different protected attributes. Similarly, Deshpande et al. [127] tackle bias by reweighting term frequency–inverse document frequency values based on fairness metrics when matching resumes with job descriptions. Burke et al. [128] introduce a mitigation method through candidate ranking. While these techniques represent an initial effort in bias mitigation, non-textual features within screening models, such as images (commonly found in online profiles), and key non-protected attributes like educational background may still contribute to perpetuating bias. This issue, where non-textual features such as visual and auditory cues contribute to bias, is particularly prevalent in the screening of candidate video or phone interviews.
3.3 Interviewing Candidates
@@
The third stage in the recruitment process involves interviewing candidates to assess their KSAOs, work values, personality, and other traits in detail [47, 91]. Over the past decade, the recruitment process has increasingly relied on AI to evaluate these traits during interviews. Various AI modalities have been utilized, including chatbot systems for candidate interaction [151, 152, 153], automated video and phone interviews [123, 36, 154], and online question-answer assessments [123, 168, 23, 150]. LLMs have also been applied for mock interviews and scoring interview answers [155]. Similarly, language models have been used to prepare job interview questions based on resume content [156] and refine pre-defined questions [7]. Commercial platforms like HireVue [123], Micro1 [110], and BrightHire [101] also heavily depend on AI to evaluate video interviews, scoring candidates based on personality traits and their fit for the job.
However, video- and image-driven AI interviewing systems have raised substantial fairness concerns, leading to both public and regulatory scrutiny, as highlighted in previous studies [140, 141, 142]. For instance, HireVue discontinued its video-based candidate scoring systems in 2021, citing concerns over transparency and potential biases in AI decision-making [36]. Illinois enacted the Artificial Intelligence Video Interview Act in 2020, becoming the first jurisdiction to regulate AI used in candidate interviews, requiring employers to disclose AI use, describe evaluation criteria, and obtain applicant consent before interviews [143]. Biases in these AI systems typically originate from training datasets containing implicit biases or unrepresentative samples, as observed in research into both video-interviewing and chatbot systems [144, 145]. For example, the widely-used First Impressions dataset, employed in the 2017 ChaLearn challenge at CVPR, was later identified to contain labels biased against participants based on gender and race [146, 147]. Such biases can be amplified through model predictions, potentially disadvantaging marginalized groups. Furthermore, automated personality trait assessments based on self-reported data have been found to lack reliability and validity, raising caution about their use in hiring decisions [157].
Similar biases have been identified in other AI-driven interview modalities, such as automatic speech recognition (ASR) systems used to transcribe candidate responses during interviews. For instance, recent research has revealed substantial accuracy disparities in ASR models when evaluating speech from non-native English speakers, people of color [154], individuals who stutter [46], and those with speech disabilities such as dysarthria [158]. Text-based interview evaluation models relying on LLMs like GPT-4 have also demonstrated biased scoring of candidate responses across demographic groups, potentially disadvantaging certain candidates based on gender or race [159]. Furthermore, research on asynchronous video interviews has shown that altering the perceived demographics of interviewees can influence evaluations, undermining fairness perceptions of the interview process [160].
To mitigate these biases in AI-based candidate evaluation, researchers have proposed various techniques. For example, Yan et al. introduced adversarial learning approaches specifically aimed at reducing demographic bias in automated personality trait predictions during interviews [147]. Booth et al. explored bias measurement techniques within affective computing systems employed for automated analysis of video-based candidate responses [149]. Additionally, significant research efforts have been directed toward improving fairness in chatbot interviewing systems by developing fairness-aware classifiers capable of detecting and mitigating harmful or biased textual interactions [148].
3.4 Selection and Evaluation
@@
The final stage of the recruitment process involves selecting one or more job candidates, followed by negotiating an offer, which is then used to evaluate the organization’s overall recruitment process [47, 91]. AI has been incorporated in several ways to assist not only in selection but also in the negotiation of job offers. For example, AI is used during job offer negotiations to provide salary and training recommendations, thereby aiding in the onboarding process of newly hired employees [168]. This is exemplified by software like Oracle Recruiting Cloud [163] which not only forecasts the probability of a job candidate accepting an offer but also allows employers to experiment with various salary structures, benefits packages, and other incentives to enhance the likelihood of acceptance [168]. Additionally, Pactum AI [161] automates various negotiation processes, such as salary negotiations and contract discussions, leveraging a multitude of factors, including job market outlook. Related studies relying upon LLMs [164, 165] and neural networks [167] have been published for forecasting job market demand and salary. Similarly, Sharma et al. [166] proposed methods to forecast employee performance, work hours, and adjust salaries accordingly.
@@
Although such AI seeks to create a fairer negotiation process by eliminating human biases, other forms of bias can still occur. For instance, a 2019 UpTurn report highlighted that Oracle and other AI-driven negotiation tools might reinforce asymmetric information between the job candidate and employer, potentially leading to an unfair process for the candidate [168]. Additionally, these tools may conflict with legislation prohibiting the use of salary history in hiring decisions [168].
@@
Job candidates who are not hired often face a lack of transparency regarding the AI-driven decision-making processes of employers. According to a survey conducted by The Talent Board, 53.5% of candidates reported receiving no feedback after the screening and interview stages, while 69.7% received no feedback after being turned down [177]. Although employers are not legally obligated to provide explanations [177], the absence of feedback can negatively impact the perceived fairness and trustworthiness of an AI-based recruitment process. Transparency is crucial at every stage of the process, including communicating the job outcome, as it enables candidates to learn and improve for future interviews. Previous studies have demonstrated that providing explanations for organizational decisions can significantly enhance trust [54, 47]. Recently, many approaches, as further described in Section 4.2, have prioritized model interpretability and explainability to enhance transparency in AI-based decision-making, a principle that can be extended to AI-driven hiring decisions. The approach by Ji et al. [167] propose a neural network for explainability of salary predictions.
@@
After finalizing hiring decisions, organizations evaluate their recruitment process based on three key criteria: validity, utility, and fairness. Validity ensures that assessments and interviews accurately reflect job KSAOs and predict job performance [47]. For instance, organizations often establish a criterion-referenced cut score as a cut-off for employee selection, based on the desired level of performance for a new employee [47]. Compliance with government guidelines, as described in Section 2.2, is also essential to validity [47]. Next, utility assesses the cost-benefit ratio of staffing, which can be influenced by earlier recruitment stages such as employee outreach and job marketing. Finally, the fairness criteria is evaluated using employee selection guidelines described in Section 2.1. However, as AI becomes more integrated into each stage of the process, it is necessary to incorporate bias detection and mitigation methods where guidelines may not adequately address algorithmic bias. For example, the uniform guidelines on employee selection could be substantially affected by the sample size of different groups. Work by Peng et al. [162] demonstrates how AI-human, or hybrid, decisions where AI assists humans, can still carry bias. Moreover, legislative definitions of fairness and employee selection guidelines may not be sufficiently detailed or up-to-date to address recent technological advancements.
4 Bias Detection and Mitigation